Derive the reduction formula


integeration(x2+a2)n/2dx = x(x2+a2)n/2÷n+1+ na2/n+1×integeration (x+a2)n/2-1dx

Use the formula to integrate integration(x2+a2)5/2

1
Expert's answer
2021-06-06T15:41:28-0400
"\int udv=uv-\int vdu"

"u=(x^2+a^2)^{n\/2}, du=nx(x^2+a^2)^{n\/2-1}dx"

"dv=dx, v=x"


"\int (x^2+a^2)^{n/2}dx=x(x^2+a^2)^{n/2}-\int x^2n(x^2+a^2)^{n/2-1}dx"

"=x(x^2+a^2)^{n/2}-n\int (x^2+a^2)^{n/2}dx+na^2\int (x^2+a^2)^{n/2-1}dx"

Then


"\int (x^2+a^2)^{n/2}dx=\dfrac{1}{n+1}x(x^2+a^2)^{n/2}"

"+\dfrac{na^2}{n+1}\int(x^2+a^2)^{n/2-1}dx, n\not=-1"

"n=5"


"\int (x^2+a^2)^{5/2}dx=\dfrac{1}{6}x(x^2+a^2)^{5/2}"

"+\dfrac{5a^2}{6}\int(x^2+a^2)^{3/2}dx"

"\int(x^2+a^2)^{3/2}dx=\dfrac{1}{4}x(x^2+a^2)^{1/2}"

"+\dfrac{3a^2}{4}\int(x^2+a^2)^{1/2}dx"

"\int(x^2+a^2)^{1/2}dx=\dfrac{1}{2}x(x^2+a^2)^{1/2}"

"+\dfrac{a^2}{2}\int(x^2+a^2)^{-1/2}dx"

"\int(x^2+a^2)^{-1/2}dx=\ln(|a\sqrt{x^2+a^2}+|x||a||)+C_1"

"\int(x^2+a^2)^{1/2}dx=\dfrac{1}{2}x(x^2+a^2)^{1/2}"

"+\dfrac{a^2}{2}\ln(|a\sqrt{x^2+a^2}+|x||a||)+C_2"